# Chaos Engineering: A Multi-Vocal Literature Review

### v1

@@ 7. Discussion
In this section, we summarize the key findings of our review and discuss their implications. We also outline open issues in chaos engineering research and discuss potential threats to our study’s validity.
@@ 7.1. Implications of Findings
  • Need for a Unified Definition: The literature reveals inconsistencies in defining chaos engineering. Although tech giants like Google, Amazon, and Netflix provide definitions that emphasize resilience testing through controlled failure injection, there is variation in focus across other companies, with some prioritizing production-level experimentation, self-healing capabilities, or anti-fragility. A unified definition can align research efforts and support the systematic adoption of chaos engineering practices, strengthening reliability across complex Information Technology (IT) environments (Siswanto et al., 2024; Jayakody and Wijayanayake, 2024; Jones, 2019).
  • Complementary Role to Traditional Testing: Chaos engineering should be viewed as a complement to traditional testing rather than a replacement (Bocetta, 2019; Palani and Gupta, 2023). While conventional testing may miss subtle vulnerabilities within complex systems, chaos engineering exposes these issues through proactive experimentation. Recognizing chaos engineering as an additional layer to traditional methods emphasizes its role in enhancing system resilience, especially in high-stakes, mission-critical environments. Integrating chaos engineering into testing workflows is encouraged for those aiming to build more robust and reliable systems.
  • Growing Alignment of Research Theory and Practice: The prevalence of solution-oriented papers in academic literature alongside experience reports in grey literature reflects an increasing alignment between research and practice in chaos engineering. This synergy allows researchers to refine models based on real-world applications while practitioners benefit from research-backed frameworks, thus bridging the gap between theory and implementation. As suggested in previous studies (Jernberg et al., 2020; Rosenthal and Jones, 2020), fostering industry-academia collaborations could further integrate academic rigor with practical insights, advancing chaos engineering as a more resilient and impactful field.
  • Targeted Tool Selection: The study suggests that organizing chaos engineering tools by factors like execution environment and automation level might help practitioners make informed decisions when selecting tools. This taxonomy enables organizations to choose tools that match their specific infrastructure, operational needs, and risk tolerance, minimizing disruptions, managing costs, and enhancing resilience testing across development stages. This targeted approach allows tools to align with specific testing phases, from initial experimentation to production-level assessments, ensuring that chaos engineering efforts are both cost-efficient and impactful, ultimately supporting system reliability.
@@
  • Standardized Practices: The findings highlight the importance of a standardized approach in chaos engineering to ensure consistent and practical application across diverse environments. Following best practices, which contribute to reliable and predictable outcomes (Kumara et al., 2021), enables organizations to build resilience and systematically manage risks. These practices, such as starting with small, controlled experiments, setting clear objectives, and automating processes where possible, support organizations in achieving reliable outcomes crucial for managing complex systems. Conversely, neglecting these practices, such as misaligned experiment design or insufficient safety mechanisms, can lead to inconsistent testing and increased system vulnerability. These findings underscore the need to balance adherence to best practices with avoiding common pitfalls, ensuring that chaos engineering aligns with broader operational goals to foster stability and reliability.
@@
  • Insights through Quantitative and Qualitative Fusion: An essential implication of the findings is that combining quantitative and qualitative evaluation methods provides a comprehensive framework for assessing and improving system stability in chaos engineering. Quantitative metrics offer objective data on performance under stress, such as error rates, MTTR, and availability. At the same time, qualitative insights, including user experience and team feedback, reveal the broader impacts of system failures. This combined approach helps organizations identify technical vulnerabilities and understand their user and business impacts, enabling targeted investments in both stability and operational efficiency to align improvements with business goals and user expectations.
@@ 7.2. Open Research Issues and Practical Challenges in Chaos Engineering
Our review identifies several unresolved issues in chaos engineering, including research gaps and organizational challenges. We examine these issues from three perspectives: people, process, and technology.
7.2.1. People
In chaos engineering, people include engineers, managers, and other team members who collaborate to design and run experiments, ensuring that the systems are thoroughly tested (Moon, 2022; Nikolaidis et al., 2021; Graglia, 2024). First, some organizations struggle with poor communication and collaboration between teams, leading to incomplete or ineffective experiments  (Zhang et al., 2019). Encouraging more vital teamwork and shared ownership could improve results by fostering better department communication (Gill, 2021; Katirtzis, 2022; Kostic, 2024). Another challenge is the shortage of specialists, making it difficult for organizations to carry out chaos experiments. This gap could be addressed by creating comprehensive training programs and certifications that allow staff to develop the necessary expertise (Dedousis et al., 2023; Nikolaidis et al., 2021; Yu et al., 2021). Finally, cultural resistance is expected, as fear of failure and blame often hinder wider adoption  (Simonsson et al., 2021; Rosenthal and Jones, 2020). Implementing cultural change initiatives that promote learning from controlled failures and reduce fear of backlash can create a more open environment for experimentation (Green, 2023).
7.2.2. Process
The process aspect refers to the specific steps and methodologies, such as planning, testing, and reporting, that guide the implementation of chaos experiments (Bairyev, 2023; Kostic, 2024; Olmstead, 2023). First, the lack of standardized guidelines across industries makes consistent application difficult (Nikolaidis et al., 2021; Cotroneo et al., 2022; Kostic, 2024). Developing a set of best practices and frameworks would help organizations measure progress and improve the effectiveness of these practices (Shah, 2021; Bremmers, 2021). The research shows that many organizations struggle to prioritize chaos experiments alongside daily operations, often finding it difficult to allocate time and resources while maintaining operational stability. Establishing transparent processes for scheduling and running these experiments during non-critical periods or as part of routine maintenance can help mitigate the issue (Kesim et al., 2020).
7.2.3. Technology
Technology refers to the systems and infrastructure used to design, execute, and monitor failure scenarios, supporting and amplifying human efforts and processes (Graglia, 2024; Olmstead, 2023). First, chaos experiments are resource-intensive and require significant computational power and storage. This makes it difficult for organizations to allocate infrastructure without affecting critical operations  (Dedousis et al., 2023; Nikolaidis et al., 2021). To manage this, companies can schedule experiments during low-traffic periods or use scalable cloud resources to reduce strain  (Kostic, 2024; Gill, 2021). Second, deploying chaos tools requires specialized knowledge, making it challenging to scale experiments consistently (Zhang et al., 2019; Yu et al., 2021). This challenge can be alleviated with better documentation, training, and platforms that simplify deployment for teams with less expertise (Bairyev, 2023). Third, many chaos engineering tools lack automation, making setup and execution time-consuming and prone to errors. Automating processes like fault injection and monitoring can reduce the burden on engineers, minimize mistakes, and improve scalability. Organizations can address this through commercial tools or internal solutions (Bairyev, 2023; Cotroneo et al., 2022; Shah, 2021). Finally, while Artificial Intelligence (AI) holds promise for further streamlining experiment design and execution, research in this area is still limited. Using AI could reduce errors, improve efficiency, and help organizations scale their resilience efforts more effectively (Zhang et al., 2023).
@@ 7.3. Threats to Validity
We considered several potential threats to the external, construct, internal, and conclusion validity (Wohlin et al., 2012) that may impact our study.
@@ 7.3.1. Threats to External Validity
External validity refers to the extent to which the study results are applicable in a broader context (Wohlin et al., 2012). Since our multivocal review incorporated studies from various online sources, our findings may have limited applicability across the wider landscape of practices and general disciplines in chaos engineering, thus posing a threat to external validity. To enhance the external validity of our findings, we conducted multiple feedback sessions throughout the analysis of the selected studies. The discussions of each session were analyzed, providing qualitative insights that allowed us to refine our research methods and the applicability of our results. Furthermore, we compiled an online appendix with all artifacts generated during our analysis, including a complete list of sources, codes, and groups (see Section  2), making them publicly accessible for further examination. Another potential threat is the risk of missing relevant literature (grey and scientific) due to alternative terminology for related concepts that may not match our search terms. We explicitly included relevant associated synonyms and terms in our search strings to address this threat. We also leveraged search engine features that inherently consider variations of terms in a search string. The items identified were rigorously assessed based on multiple quality dimensions (Wohlin, 2014). Finally, while there are numerous chaos engineering tools, we limited our analysis to ten due to practicality. To mitigate this limitation, we selected the ten most relevant tools currently used by practitioners. A detailed comparison of these chosen tools may provide additional insight, which we propose for future research.
@@ 7.3.2. Threats to Construct and Internal Validity
Construct validity pertains to the generalizability of the study’s constructs, while internal validity relates to the credibility of the methods used for data analysis, including possible biases (Wohlin et al., 2012). To reduce potential threats to construct and internal validity, we implemented multiple triangulation rounds, inter-rater reliability assessments, and quality control measures (see Section 2). These methods helped minimize potential biases, such as observer and interpretive biases, thereby enhancing the validity of our data analysis. Additional studies and codes were incorporated, although no further inter-rater evaluation was conducted. In particular, some risk of observer bias remains inherent in this methodology (Kumara et al., 2021).
@@ 7.3.3. Threats to Conclusion Validity
Conclusion validity refers to the extent to which the conclusions of the study are credibly supported by the available data  (Wohlin et al., 2012). To address this threat, we used thematic coding and inter-rater reliability assessments to reduce observer and interpretive biases, aiming to analyze the data thoroughly. Furthermore, the study conclusions were independently derived by different authors and subsequently reviewed and validated against the primary studies selected during a feedback session.
8. Related Surveys
In this section, we consider key surveys on resilience, robustness, and fault injection techniques, exploring their relationship to the proactive focus of chaos engineering. For example, Natella et al. conducted an extensive survey on Software Fault Injection (SFI) techniques, categorizing methods by fault types, injection approaches, and application domains. These techniques are valuable for assessing software reliability in controlled settings. However, SFI techniques are generally applied in pre-production or isolated environments rather than in the live, iterative fault-testing scenarios central to chaos engineering (Natella et al., 2016). Another study, conducted by Colman-Meixner et al. (Colman-Meixner et al., 2016), explores resilience strategies across cloud infrastructure layers, such as physical servers, virtualization, and application components. Their work proposes a multi-layered framework with redundancy and failover mechanisms to ensure continuity during failures. While this approach aligns with the resilience goals of chaos engineering, the focus remains on structured recovery strategies within controlled environments rather than proactive fault injection in live settings. Laranjeiro et al. (Laranjeiro et al., 2021) reviewed robustness assessment techniques for systems such as operating systems and middleware and discussed methods such as fault injection and API testing to evaluate dependability under stress. These approaches aim to identify weaknesses but emphasize static testing, which does not include the proactive, live-system experiments that chaos engineering relies on. More recently, Rouholamini et al. (Rouholamini et al., 2020) presented a study on proactive self-healing techniques in cloud computing, categorizing strategies such as automated fault detection, anomaly recognition, and self-correction mechanisms. These techniques improve system resilience by addressing faults as they arise, but differ from the chaos engineering approach of using intentional fault injection to reveal latent vulnerabilities preemptively.
In general, there is a significant knowledge gap in the area of proactive and iterative resilience testing directly within live production environments. Moreover, there are no comprehensive reviews of the literature on chaos engineering. Therefore, this review paper systematically analyzes, assesses, and summarizes the existing literature to clarify the unique contributions of chaos engineering to the domain of system resilience.
@@ 9. Conclusion and Future Work
As modern distributed systems become increasingly complex, chaos engineering offers a proactive approach to ensure stability by introducing controlled failures that reveal system weaknesses, prevent outages, and support high performance under stress. This study bridges the gap between research insight and industry practices through a multivocal literature review (MLR) of 88 sources from academic and grey literature. To promote clarity across academia and industry, we proposed a unified definition of chaos engineering, reinforcing its role as a vital complement to traditional testing. We identified the main functionalities of a chaos engineering platform and compiled a set of principles that can guide systematically planning and conducting chaos experiments. In addition, we presented the key components of a chaos engineering platform and identified 12 quality requirements to guide the selection and evaluation of the platform. We also identified eight qualitative and four quantitative metrics essential for evaluating the impact of chaos experiments, ensuring that chaos engineering meaningfully enhances system robustness. Our review highlights four technical and three socio-technical challenges that chaos engineering seeks to resolve. To support organizations in structuring chaos engineering practices, we developed a taxonomy that organizes tools and techniques by environment, automation mode, automation strategy, and deployment stage while identifying the ten most commonly used tools within this framework. This categorization is valuable to practitioners and researchers when selecting tools that align with specific infrastructure needs and risk profiles. In addition, we identified six best practices and five common pitfalls in the implementation of chaos engineering. These practices and potential pitfalls guide organizations in optimizing their chaos engineering initiatives to foster safer and more robust systems. Our findings also reveal that solution-oriented research in academia and practical experience in industry converge, underscoring chaos engineering’s essential role in strengthening modern distributed systems. Finally, we outlined several research and practical challenges that must be addressed to realize chaos engineering’s vision entirely. We believe that these identified open issues can guide future research in the domain of chaos engineering.
@@
In future work, we plan to use GitHub repository mining to identify various open-source projects that implement chaos engineering practices and utilize associated tools. This exploration will complement the ten tools analyzed in this literature review, offering broader insights into the prevalence and practical application of chaos engineering in diverse projects. Additionally, we aim to apply chaos engineering to improve the robustness of AI-enabled systems by evaluating how such systems handle disruptions in accuracy, adaptability, and consistency, ensuring that they continue to deliver reliable performance even under challenging conditions.

### v2

@@ 8. Discussion
In this section, we summarize the key findings of our review and discuss their implications across RQ1 to RQ4. We also outline open issues in chaos engineering research and discuss potential threats to the validity of our study.
@@ 8.1. Implications of Findings
  • Extending Reliability Testing through Chaos Engineering: Chaos engineering should be understood not as a replacement for established reliability or software testing techniques, but as an extension that focuses on system behavior under failure and uncertainty. While traditional methods, such as unit tests, integration tests, or stress tests, validate expected behaviors in controlled environments, chaos engineering introduces carefully scoped disruptions to examine how systems respond and recover from unexpected faults. This approach contributes a complementary dimension to reliability assessment, particularly in distributed or production-like environments where emergent failures often go undetected. Rather than displacing existing practices, chaos engineering enhances them by validating resilience under real-world operational stress.
  • Targeted Tool Selection (RQ3.1): This study introduces a taxonomy that helps practitioners select chaos engineering tools by considering factors such as execution environment, automation mode, and deployment stage (see Table 13). The taxonomy is intended to support informed decision-making by aligning tool capabilities with the specific infrastructure and operational needs of an organization. In practice, a practitioner would begin by assessing the architectural characteristics of their system, such as whether it runs on Kubernetes, serverless platforms, or hybrid cloud, and then identify tools from the taxonomy that match these conditions and the intended stage of testing, from early experimentation to full production deployment. Applying the taxonomy effectively also requires careful consideration of how well a tool integrates with existing workflows and team expertise. Practitioners must ensure that selected tools support their resilience objectives without introducing unnecessary complexity or risk. For example, using a highly automated tool in an immature testing environment could result in poorly scoped experiments or overlooked failure modes. Likewise, choosing tools that lack compatibility with the target infrastructure may lead to inconclusive or misleading outcomes. By mapping tools to their most appropriate use cases, the taxonomy can help mitigate such risks and promote more efficient and targeted adoption of chaos engineering practices across diverse technical landscapes.
@@
  • Standardized Practices (RQ3.2): The findings highlight the importance of a standardized approach in chaos engineering to ensure consistent and practical application across diverse environments. Following best practices, which contribute to reliable and predictable resilience outcomes, such as faster recovery times and minimal service disruption (Kumara et al., 2021), enables organizations to build resilience and systematically manage risks. These practices, such as starting with small, controlled experiments, setting clear objectives, and automating processes where possible, support organizations in achieving reliable outcomes crucial for managing complex systems (see Table 14). Conversely, neglecting these practices, such as misaligned experiment design or insufficient safety mechanisms (see Table 15), can lead to inconsistent testing and increased system vulnerability. These findings underscore the need to balance adherence to best practices with avoiding common pitfalls, ensuring that chaos engineering aligns with broader operational goals to foster stability and reliability.
@@
  • Insights through Quantitative and Qualitative Fusion (RQ3.3): Another key implication of our findings is that combining quantitative and qualitative evaluation methods enhances the assessment and improvement of system stability in chaos engineering. Quantitative metrics provide objective data on performance under stress, including error rates, mean time to resolution, and availability. At the same time, qualitative insights, including user experience and team feedback, reveal the broader impacts of system failures. This combined approach enables organizations to identify technical vulnerabilities and understand their user and business impacts, allowing for targeted investments in both system stability and operational efficiency. This alignment of improvements with business goals and user expectations facilitates the development of solutions that effectively enhance system resilience, reduce failure impact, and improve user experience.
  • Growing Alignment of Research Theory and Practice (RQ4): Our review empirically demonstrates the growing alignment between academic research and industry practice in chaos engineering. By classifying 96 primary sources, including both peer-reviewed papers and grey literature, we observed that solution proposal papers dominated academic publications between 2016 and 2023, while grey literature during the same period showed a slight increase in experience reports (see Figure 11). This dual trend reveals a convergence between theoretical contributions and practitioner-led experimentation, further reflected in the types of publication venues: top-ranked academic venues published novel frameworks and techniques, while widely cited blog posts detailed real-world implementation case studies. By capturing both perspectives, our study illustrates how academic research and industry practices are increasingly aligned, with researchers drawing insights from real-world applications and practitioners adopting concepts that resonate with emerging scholarly work to guide system resilience strategies. This evolving interplay signals an opportunity for deeper collaboration across academia and industry to accelerate innovation in chaos experiment design, such as advancements in automated fault orchestration, address implementation gaps, such as the lack of robust rollback or abort mechanisms in several research prototypes, and broaden the real-world adoption of chaos engineering.
@@ 8.2. Open Research Issues and Practical Challenges in Chaos Engineering
Our review highlights several ongoing challenges that limit the adoption and scalability of chaos engineering. These include cultural resistance, skill gaps, and resource overhead. The following subsections outline three key themes based on current literature and practice.
8.2.1. Organizational and Cultural Challenges
The findings highlight that organizational resistance, risk aversion, and a blame culture can pose significant barriers to the adoption of chaos engineering (Krishnamurthy, 2021; Bocetta, 2019). Although chaos engineering promotes resilience, many teams perceive fault injection, especially in production, as disruptive and risky (Sbiai et al., 2023; Zhang et al., 2019; Dedousis et al., 2023). This perception discourages widespread adoption, as teams are often reluctant to participate in experiments that could disrupt production or damage service reliability (Krishnamurthy, 2021; Ji et al., 2023). Adopting chaos engineering requires a mindset shift: from protecting systems at all costs to embracing failure as a tool for improvement (Bocetta, 2019; Bremmers, 2021). However, this shift is rarely supported by leadership and is often undermined by siloed teams, poor communication, and inconsistent reliability practices (Katirtzis, 2022; Dua, 2024; Abdul, 2024). Without cultural alignment and collaboration, chaos engineering remains confined to isolated efforts with limited organizational impact (Sbiai et al., 2023; Tivey et al., 2019). While these challenges are frequently noted in industry reports, academic research has yet to explore them systematically. There is a need for empirical studies that examine how factors like culture, team structure, and leadership influence chaos engineering adoption. Research could also contribute by developing and validating readiness frameworks or maturity models to guide organizations in scaling these practices effectively.
8.2.2. Skill and Expertise Gaps
Chaos Engineering requires specialized skills that many teams currently lack, creating a major barrier to adoption (Torkura et al., 2020; Kostic, 2024). First, effective implementation depends on understanding distributed systems, runtime behavior, and formal models like metric temporal logic, areas in which most practitioners have limited training (Frank et al., 2023; Nikolaidis et al., 2021). Next, managing experiment configurations, such as complex YAML files with interdependent settings, demands precision and familiarity with system internals (Kassab et al., 2022). Furthermore, many chaos engineering practices rely on intuition rather than standardized procedures, leading to inconsistent outcomes and reduced reliability (Nikolaidis et al., 2021; Vinisky, 2024). Adding to this challenge, some experiments require interaction with low-level components, such as kernel modules or orchestration layers, which increases the technical entry barrier (Katirtzis, 2022; Mooney, 2023). Although these challenges are recognized in both academic and practitioner sources, there is limited empirical research on how teams acquire and apply these skills in practice. Future work should focus on identifying specific knowledge gaps, evaluating effective training strategies, and designing practical tools and reusable templates to support broader and more consistent adoption (Bairyev, 2023; Cotroneo et al., 2022).
8.2.3. Resource and Operational Constraints
Chaos engineering demands significant time, computing resources, expertise, and cost, which often limits routine adoption  (Vu et al., 2022; Poltronieri et al., 2022; Zhang et al., 2021a). Setting up environments, managing telemetry, and iterating experiments require sustained effort, especially for smaller or time-constrained teams (Gill, 2021; Zhang et al., 2021a). Deploying testbeds can take considerable time and need expert oversight (Kassab et al., 2022). These challenges scale with system size and are intensified by limited infrastructure, cloud expertise, and the financial overhead of tooling and operations (Katirtzis, 2022; Durai, 2022; Kostic, 2024). As systems become increasingly complex, chaos engineering becomes more challenging to scale due to service heterogeneity and coordination overhead. Security and data concerns also arise. For example, experiments may expose vulnerabilities and generate sensitive telemetry. Managing this data poses storage, privacy, and analysis challenges. These factors—including time, complexity, and cost—often clash with delivery timelines, leading to chaos engineering being deprioritized despite its long-term value (Krivas and Portela, 2020; Kostic, 2024; Bairyev, 2023). Although frequently reported in practice, there is little empirical research on how teams manage these constraints. Future work could investigate how organizations reduce the setup overhead of test environments or prioritize experiments to align with available resources. Such studies would help tailor chaos engineering to practical operational limits.
@@ 8.3. Threats to Validity
We considered potential threats to the external, construct, internal, and conclusion validity (Wohlin et al., 2012) that may impact our study.
@@ 8.3.1. Threats to External Validity
External validity refers to the extent to which our findings can be generalized beyond the specific context of our study (Wohlin et al., 2012). While this multivocal literature review (MLR) included both academic and grey sources, there is a possibility that some perspectives or practices in chaos engineering may not have been captured. A particular threat arises from the risk of excluding relevant literature that uses alternative terminology, such as resilience testing or fault injection, rather than chaos engineering. While we crafted inclusive search strings using multiple synonyms and related terms (see Table 3), some relevant works may still have been omitted. This limitation is common in MLRs and is explicitly acknowledged. Future studies may improve coverage by expanding search terms to encompass other reliability testing practices. Another limitation stems from the taxonomy itself, which has not yet undergone formal external validation. Although developed through thematic analysis and informed by structured internal review sessions, it remains untested in practical or industry settings. As such, its generalizability should be interpreted with caution until further validation studies are conducted. Finally, our tool analysis focused on ten widely adopted chaos engineering tools due to feasibility constraints. While these tools represent current practitioner usage, broader tool coverage, including lesser-known or homegrown solutions, could yield deeper insights. We consider this an avenue for future work.
@@ 8.3.2. Threats to Construct and Internal Validity
Construct validity concerns whether the study accurately captures the concepts it aims to investigate, while internal validity refers to the soundness of the methods used and the control over potential biases (Wohlin et al., 2012). To address these concerns, we conducted at least four structured feedback sessions during our analysis. Discussions from these sessions were qualitatively examined and used to fine-tune both our methodological choices and the relevance of our findings. We have also compiled an online appendix (see Section 3.6) that contains all key artifacts, including the full list of sources, extracted codes, thematic groupings, and taxonomy classifications, to ensure transparency and reproducibility. We also adopted triangulation by synthesizing both academic and grey literature sources, thereby reducing our reliance on a single perspective. Inter-rater reliability was formally assessed during the study selection phase, using Cohen’s Kappa, which indicated strong agreement for both academic and grey literature. This ensured consistency in applying inclusion and exclusion criteria. The first author conducted manual coding, which was independently reviewed by the second author. Disagreements were resolved collaboratively through discussion. This collaborative process was applied throughout all coding phases to ensure consistency and reduce interpretation bias. Despite these precautions, some risk of observer bias remains inherent in qualitative research. However, we believe that our triangulation of sources, structured collaboration, and transparent artifact release significantly mitigate this threat.
@@ 8.3.3. Threats to Conclusion Validity
Threats to conclusion validity concern the degree to which the study’s conclusions are reasonably based on the data analyzed (Wohlin et al., 2012). To mitigate this, we applied thematic coding using ATLAS.ti tool, and employed a structured analysis process to reduce observer and interpretation biases. Coding and classification were guided by a consensus-driven protocol between the first and second authors, ensuring consistency in the derivation of findings. Both authors independently formulated the conclusions presented in this study and later validated them through collaborative discussion. These conclusions were cross-checked against the original source material to ensure they remained grounded in the data. While our synthesis integrated both academic and grey literature, we acknowledge that our interpretations are limited to the scope of the selected articles. In particular, the absence of external expert validation for the taxonomy remains a limitation. Future research may benefit from follow-up evaluations, broader empirical testing, or practitioner feedback to confirm the applicability of our findings in real-world settings. We also acknowledge that our results are limited to the quality of the grey literature and peer-reviewed publications, which we mitigate by conducting a quality assessment analysis.
@@ 9. Conclusion and Future Work
As modern distributed systems become increasingly complex, chaos engineering offers a proactive approach to ensure stability by introducing controlled failures that reveal system weaknesses, prevent outages, and support high performance under stress. This study bridges the gap between research insight and industry practices through a multivocal literature review (MLR) of 96 sources from academic and grey literature. To promote clarity across academia and industry, we proposed a unified definition of chaos engineering, reinforcing its role as a vital complement to traditional testing. We identified the main functionalities of a chaos engineering platform and compiled a set of activities that can guide systematic planning and conducting chaos experiments. In addition, we presented the key components of a chaos engineering platform and identified eleven quality requirements to guide the selection and evaluation of the platform. We also identified eight qualitative and four quantitative metrics essential for evaluating the impact of chaos experiments, ensuring that chaos engineering meaningfully enhances system robustness. Our review highlights three technical and four socio-technical challenges that chaos engineering seeks to resolve. To support organizations in structuring chaos engineering practices, we developed a taxonomy that organizes tools and techniques by environment, automation mode, automation strategy, and deployment stage while identifying the ten most commonly used tools within this framework. This categorization is valuable to practitioners and researchers when selecting tools that align with specific infrastructure needs and risk profiles. In addition, we identified six best practices and five common pitfalls in the implementation of chaos engineering. These practices and potential pitfalls guide organizations in optimizing their chaos engineering initiatives to foster safer and more robust systems. Our findings also reveal that solution-oriented research in academia and practical experience in industry converge, underscoring the essential role of chaos engineering in strengthening modern distributed systems. Finally, we outlined several research and practical challenges that must be addressed to realize the vision of chaos engineering fully. We believe that these identified open issues can guide future research in the domain of chaos engineering.
@@
In future work, we plan to use GitHub repository mining to identify various open-source projects that implement chaos engineering practices and utilize associated tools. This exploration will complement the ten tools analyzed in this literature review, offering broader insights into the prevalence and practical application of chaos engineering in diverse projects. Additionally, we aim to apply chaos engineering to enhance the robustness of AI-enabled systems by evaluating how these systems respond to disruptions in accuracy, adaptability, and consistency, ensuring they continue to deliver reliable performance even under challenging conditions.